OpenStudy (anonymous):

x over a - y over b =1 solve for y

OpenStudy (anonymous):

y=xb/a-b

OpenStudy (anonymous):

can you show the steps?

OpenStudy (anonymous):

[x/a-y/b=1] [x(b)-y(a)/(ab)=1] [x(b)-y(a)=ab] [-y(a)= ab-xb] [-y=(ab-xb)/a] [y=-[(ab-xb)/a]]

OpenStudy (anonymous):

not sure thats right still thinking

OpenStudy (anonymous):

thats a note for myself

OpenStudy (anonymous):

i would think that would work but i could be wrong

OpenStudy (anonymous):

x/a-y/b=1 x(b)-y(a)/(ab)=1 x(b)-y(a)=ab -y(a)= ab-xb -y=(ab-xb)/a y=-(ab-xb)/a

OpenStudy (anonymous):

do you have the answer?

OpenStudy (anonymous):

im good at finding solutions when i have the answer lol